OpenStudy (anonymous):

solve for x 8-11x=5+4/3(2-x) 4/3 is a fraction read 4 3rds

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

when solving this, x becomes a decimal point not a whole number. is that what you are asked for?

OpenStudy (anonymous):

i just need to solve for x i know it will be a fraction, decimal, or whatever. This problem is driving me nuts!

OpenStudy (anonymous):

i got \[x=\frac{-1}{37}\]

OpenStudy (anonymous):

steps??

OpenStudy (anonymous):

Is the equation this? \[8-11x=5+\frac{4}{3}(2-x) \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The first thing I would do is multiply both sides by 3 to get rid of the fraction.

OpenStudy (anonymous):

thats where im stuck what exactly do I multiply by 3?

OpenStudy (anonymous):

Everything.

OpenStudy (anonymous):

\[8-11x=5+\frac{4}{3}(2-x)\]\[\implies 3[8-11x] = 3[5+\frac{4}{3}(2-x)]\]

OpenStudy (anonymous):

\[8-11x=5+\frac{4}{3}(2-x)\] distribute \[\frac{4}{3}\] so it equals to \[8-11x=5+\frac{8}{3}-\frac{4}{3}x\] in one step i added \[(5+\frac{4}{3}) \] and -11x from left to right we get the following \[8=\frac{23}{3}-\frac{37}{3}x\] \[subtract \frac{23}{3} \to both sides\] \[\frac{1}{3}=-\frac{37}{3}x\] then divide.

OpenStudy (anonymous):

That works too, but lots of fractions that way.

OpenStudy (anonymous):

polpak, i solved it this way. I like fractions.

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

Yo dawg. I herd you liek fractions so I put some fractions in your fractions so you can divide when you divide.

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Jwaddell, do you like fractions?

OpenStudy (anonymous):

i hate them they cause issues like this

OpenStudy (anonymous):

Jwaddell no matter what way you solve it. The answer is the same. if you want to. polpak or my self could help you solve by getting rid of fractions first.

OpenStudy (anonymous):

Ok, well if you multiply both sides of the equation you will rid yourself of the fraction.

OpenStudy (anonymous):

everyway ive tried comes out different

OpenStudy (anonymous):

multiply by 3.

OpenStudy (anonymous):

\[3[8−11x]=3[5+{4\over 3}(2−x)]\]

OpenStudy (anonymous):

So distribute the 3 on the left and you get: 24 - 33x = 15 + 4(2-x)

OpenStudy (anonymous):

got that

OpenStudy (anonymous):

Ok, so now distribute the 4 on the right hand side

OpenStudy (anonymous):

What have you got?

OpenStudy (anonymous):

24-33x=15+8-4x

OpenStudy (anonymous):

Perfect. Now combine the 15 + 8

OpenStudy (anonymous):

k

OpenStudy (anonymous):

23

OpenStudy (anonymous):

Ok so you have \[24 - 33x = 23 - 4x\] So lets add 33x to both sides.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

24=23-37x???

OpenStudy (anonymous):

Nope.. -4x + 33x = ?

OpenStudy (anonymous):

or is it -4+33 to = 29x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

k

OpenStudy (anonymous):

So we have 24 = 23 + 29x Now we subtract 23 from both sides.

OpenStudy (anonymous):

so then 1=29x

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

1/29=x

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Perhaps beehive should not like fractions so much ;p

OpenStudy (anonymous):

i guess so. I hurry a lot too much.

OpenStudy (anonymous):

thank you!!!!

OpenStudy (anonymous):

very welcome. Be sure to review what we did so you can understand how to do it again in the future.

OpenStudy (anonymous):

i will